OpenStudy (anonymous):

so if i do 18 questions on a 20 question whats my average

OpenStudy (anonymous):

What?

OpenStudy (sleepyhead314):

you mean your grade if you get all the questions correct? (correct questions) / (total) * (100) = percent of correct please ask this question in the mathematics section next time ^_^

OpenStudy (anonymous):

0.9

OpenStudy (anonymous):

That would be 90% on your grade.

OpenStudy (anonymous):

sorry didn't know was in the wrong section

OpenStudy (anonymous):

@sleepyhead314 I think he meant if he got 18/20 questions right.

OpenStudy (anonymous):

yeah

OpenStudy (sleepyhead314):

yes (correct questions) / (total) * (100) = percent of correct so 18 / 20 * 100 = 90%